# Enhancing Biomedical Knowledge Discovery for Diseases: An End-To-End Open-Source Framework

### v1

@@ 1 Introduction
Knowledge discovery (Wang et al., 2023; Shu and Ye, 2023) is a pivotal research domain due to the surge in publications, which makes keeping up with new findings challenging, necessitating automated knowledge extraction and processing. Of particular concern is the biomedical literature, where updates occur with ever-accelerating frequency (Fig. 1). Despite advances in healthcare, many diseases, such as Alzheimer’s disease (AD) (Trejo-Lopez et al., 2023; Scheltens et al., 2021) and multiple sclerosis (McGinley et al., 2021; Attfield et al., 2022), lack effective cures. Additionally, over 1,200 rare disorders have limited or no cures according to the National Organization for Rare Disorders. [^2] https://rarediseases.org/rare-diseases/ Discovering new scientific insights from research papers can expedite disease understanding and accelerate cure development.
@@
  • Two annotated datasets for RS and AD provide gold labels for semantic relations, aiding disease knowledge discovery research. [^3] The description of distantly supervised datasets for weakly supervised scenarios is included in Appendix F.
@@ 2 Data Pipeline
Mention extraction. MetaMapLite (Aronson, 2001) (open-source BSD License) is provided by the National Library of Medicine (NLM) for extracting biomedical entities and mapping them to Concept Unique Identifiers (CUIs) within UMLS. The tool is updated every two years to incorporate the latest medical terminology and to ensure its accuracy in extraction and mapping. MetaMapLite simultaneously extracts mentions and links them to UMLS in one step, efficiently associating mentions with their corresponding CUIs. We detect a diverse range of entities, spanning 82 unique semantic types and covering a broad spectrum of biomedical concepts, including diseases, biologically active substances, anatomical structures, genes, and more. Detailed entity detection often leads to overlapping or successive entities in the text. To address this, our pipeline incorporates a merging strategy that consolidates overlapping or subsequent entities into cohesive units. For example, in the sentence: "To test norepinephrine augmentation as a potential disease-modifying therapy, we performed a biomarker-driven phase II trial of atomoxetine, a clinically-approved norepinephrine transporter inhibitor, in subjects with mild cognitive impairment due to AD.", the subsequent relevant mentions norepinephrine transporter and inhibitor are merged to one entity.
@@ 2.1 Dataset Creation
The result of the expert annotation yields two curated datasets. The Relation Detection dataset for Rett Syndrome (ReDReS) contains 601 sentences with 5,259 instances and 1,148 unique CUIs (Tab. 1). The inter-annotator agreement is measured using the Fleiss kappa score (McHugh, 2012), resulting in 0.6143 in the multi-class setup (4 classes) and indicating substantial agreement among annotators (Landis and Koch, 1977). In the binary setup (relation or no relation), the Fleiss kappa score is 0.7139. The Relation Detection dataset for Alzheimer’s Disease (ReDAD) comprises 641 sentences with 8,565 instances and 1,480 unique CUIs (Tab. 1). The Fleiss kappa score is 0.6403 in the multi-class setup and 0.7064 in the binary setup, showing substantial consensus among annotators. The final labels are determined through majority voting, leveraging the labels provided by each expert. While the label distribution across classes is relatively balanced, the negative class is under-represented with 97 and 125 instances in ReDReS and ReDAD respectively (Tab. 1). Each dataset is randomly split into train, development, and test sets.
@@ 3 Models
Figure 3: Model Architecture of LaMReDA, LaMReDM (left), and LaMEL (right): Each model encodes the input sequence using PubMedBERT (large or base). For LaMReDA and LaMReDM, different tokens define the relation representation (A-P), passed through a linear projection layer, a dropout layer, and then a classification layer for prediction. The symbol # denotes element-wise addition and multiplication for LaMReDA and LaMReDM, respectively. For LaMEL, different tokens construct the entity representation (A-H), which are sent through a dropout layer and a linear layer to extract the projected entity representations.
@@ 3.1 LaMEL model
LaMEL learns an embedding space optimized for relation detection (Fig. 3). As the backbone language model (LM), we opt for PubMedBERT (Gu et al., 2021; Tinn et al., 2023) (MIT License), available in both uncased base and uncased large versions. [^8] HuggingFace’s Transformers library (Wolf et al., 2019) PubMedBERT is pretrained on the PubMed corpus, making it well-suited for our task as the curated datasets consist of sentences of abstracts from PubMed papers. Leveraging PubMedBERT ensures that the model can capture the language patterns prevalent in biomedical text. Following the LM encoding, we construct the representation of each entity by extracting its contextualized embedding $E_{i}$ corresponding to each entity $e_{i}$ from the encoded sequence. Subsequently, the entity representations are projected to the embedding space using a linear layer without changing the embedding dimension. The final prediction is based on cosine similarity between the two projected entity representations. If the cosine similarity exceeds a predefined threshold, the model predicts that there is a semantic relation between the two entities. We experiment with diverse strategies for learning entity representations (Fig. 3), aiming to optimize the effectiveness of the embedding space for the relation detection task. The explored types of entity representation $E$ are:
@@
where $\{E_{A},E_{B},E_{D},E_{E},E_{F},E_{G},E_{H}\}\in\mathbb{R}^{d}$ and $E_{C}\in\mathbb{R}^{2d}$, $d$ is the embedding size of PubMedBERT base (768) and PubMedBERT large (1024), ; defines the concatenation, * holds for the element-wise multiplication, $t_{[ent]}$, $t_{[/ent]}$ are the embeddings of the start and end special tokens of the entities, $[t_{E}]$ and $[t_{Inter}]$ are the averaged pooled representation of the entities and the intermediate tokens between the entities respectively.
@@ 3.2 LaMReD model
LaMReD provides two variations that differ in information synthesis (Fig. 3), aiming to explore the potential effect of different aggregations (Theodoropoulos and Moens, 2023). LaMReDA utilizes element-wise addition to aggregate the entities’ representations, while LaMReDM employs element-wise multiplication. The input text is encoded using PubMedBERT (base or large). Following LM encoding, we construct the relation representation by sampling and aggregating tokens from the input sequence. This step enables the model to capture essential features and contextual information relevant to semantic relation classification. To mitigate the risk of overfitting and enhance model generalization, we incorporate a dropout layer (Srivastava et al., 2014) with a probability of 0.3. The linear classification layer takes the aggregated representation and outputs the predicted label.
@@
where $\{R_{A},R_{B},R_{C},R_{D},R_{E},R_{F},R_{G},R_{H},R_{I},\\ R_{J},R_{K},R_{L},R_{M},R_{N},R_{O},R_{P}\}\in\mathbb{R}^{d}$, d is the embedding size of PubMedBERT base (768) and PubMedBERT large (1024), $f()$ is the aggregation function, element-wise addition for LaMReDA and element-wise multiplication for LaMReDM, $l()$ is a linear projection layer with dimension equal to the embedding size, $t_{[ent]_{1}}$, $t_{[/ent]_{1}}$, $t_{[ent]_{2}}$, and $t_{[/ent]_{2}}$ are the embeddings of the start and end special tokens of the first and second entity and $t_{[CLS]}$ is the representation of the special token [CLS]. We define the averaged pooled representation of the entities and the intermediate tokens between the entities as $[t_{E1}]$, $[t_{E2}]$, and $[t_{Inter}]$ correspondingly. In equations 23 and 24, we utilize the localized context vector $cv$ [^9] Additional information is provided in Appendix D. that utilizes the attention heads to locate relevant context for the entity pair and was introduced in ATLOP (Zhou et al., 2021), a state-of-the-art model in document-level relation extraction.
@@ 4 Results
Tables 2 and 3 report the F1-scores for LaMEL LaMReDA, and LaMReDM, models on the ReDReS and ReDAD datasets. Each cell (except for cross-disease experiments) displays two values: the average F1-score from 10 runs on the original test set (Tab. 1) and the average F1-score from a 5-fold cross-validation. The models perform well across all relation (A-P) and entity (A-H) representations, showing their ability to learn meaningful representations for the semantic relation task regardless of initial token selection. However, we observe patterns regarding the relation representations. In the binary setup, relation representation $R_{G}$ (Eq. 15) yields strong results for both datasets, suggesting that including the [CLS] token representation might be beneficial. In the multi-class setup, relation representations $R_{L}$ (Eq. 20), $R_{J}$ (Eq. 18), and $R_{O}$ (Eq. 23) are effective for both datasets, indicating that the surrounding context is crucial for the more complex task, as $R_{L}$ and $R_{J}$ include the averaged pooled representation of intermediate tokens between entities, and $R_{O}$ leverages the context vector (Zhou et al., 2021). The intra-model comparison reveals that over-parameterization tends to be useful. Using PubMedBERT large generally results in better performance than the base alternative. The PubMedBERT base shows superior performance mainly only in experiments using the original splits of ReDReS (Tab. 1). LaMEL is highly competitive with LaMReDA and LaMReDM, indicating that learning entity embedding spaces optimized for relation detection is promising. LaMEL achieves the highest performance in the 5-fold setup of ReDAD and the original setup of ReDReS, with F1-scores of 91.03% and 91.25% respectively.
@@
Table 2: LaMEL Results (%) in binary setup (PubMedBERT $\square$: base, $\blacksquare$:large): Each cell (unless cross-disease experiments) shows the average F1-score from 10 runs (original test set) and from 5-fold cross-validation setup.
The inter-model comparison across the same relation representations indicates that the aggregation function does not significantly impact relation detection tasks. Neither LaMReDA (element-wise addition) nor LaMReDM (element-wise multiplication) show a clear advantage over the other. This suggests that the transformer layers of PubMedBERT and the projection layer $l()$ preceding the aggregation are effectively trained in both models to encode the essential information for relation detection, regardless of the aggregation function used. The cross-disease experiments underscore the robustness of the models in both binary and multi-class setups. This robustness supports transfer learning (Zhuang et al., 2020) in semantic relation detection, extending to other diseases, highlighting the potential for broader applications and research endeavors in knowledge discovery.
@@
Table 3: LaMReDA and LaMReDM Results (%) in binary and multi-class setup (PubMedBERT $\square$:base,$\blacksquare$:large): Each cell (unless cross-disease experiments) shows the average F1-score from 10 runs (original test set) and from 5-fold cross-validation setup.
Human Performance. To assess and compare to human performance, two additional experts identify the relation type in a random sample of 300 instances from the test set of each dataset (Tab. 1). The evaluation ground truth is based on the original test set labels. In the binary setup, the average F1-score ranges from 92.14 for ReDReS to 91.87 for ReDAD. The LaMReDA, LaMReDM, and LMEL models achieve performance comparable to human experts, indicating a high ability to detect semantic relations. Multi-class macro F1-scores range from 85.23 (micro: 85.45) to 85.76 (micro: 85.87) for ReDReS and ReDAD, respectively. Compared to human experts, all models show a performance gap, highlighting that identifying more complex aspects of semantic relation is a challenging task.
Baseline performance - lower bound. We randomly assign labels based on the training data’s class distribution (Tab 1). In the binary setup, the baseline achieves F1-scores of 54% (ReDReS) and 53.16% (ReDAD). For the multi-class setup, the macro F1-scores range from 32.05% to 32.43%, stressing the task’s difficulty, particularly for distinguishing various semantic relations (multi-class) [^11] More information is available in Appendix E..
@@ 5 Probing
This study probes PubMedBERT’s ability to capture semantic relations between entities. We explore different transformer layer representations and attention scores per layer and attention head. Averaged pooled entity representations are extracted from each layer, followed by training a linear classification layer. We test relation representations $R_{D}$, $R_{O}$, and $R_{P}$ (Eq. 12, 23, 24) of LaMReDA and LaMReDM to assess the impact of the context vector. Out-of-the-box representations are evaluated without the projection linear layer $l()$. We also extract average attention scores of tokens for each entity towards the other across each layer and head, concatenating these into a feature vector for training a linear classification layer. Following Chizhikova et al. (2022), we also train the classification layer using average attention scores between the two entities across all layers.
@@
Figure 4 shows the results of probing experiments in the binary setup using ReDReS and PubMedBERT base. [^12] Additional probing experiments in Appendix G. The 10th and 11th layers provide the most informative representations for relation types ($R_{D}$, $R_{O}$, and $R_{P}$). The $R_{D}$ representation, using element-wise multiplication, outperforms other representations in intra-layer comparisons, suggesting its effectiveness without end-to-end training. However, as highlighted in section 4, inter-model comparisons indicate that the transformer layers and projection layer $l()$ capture crucial information for relation detection, regardless of the aggregation function. Using context vectors with $R_{O}$ and $R_{P}$ generally offers no advantage, though $R_{O}$ from the 10th and 11th layers performs well, indicating possibly meaningful localized context. Attention scores between entities in the 12th layer yield the best performance, surpassing the baseline of using scores from all layers, indicating strong attention between the entities in the last layer. Figure 4 reveals that the 6th and 9th attention heads are most informative for relation detection.

### v2

# Enhancing Biomedical Knowledge Discovery for Diseases: An Open-Source Framework Applied on Rett Syndrome and Alzheimer’s Disease

@@ 1 Introduction
Knowledge discovery (Wang et al., 2023; Shu and Ye, 2023) is a pivotal research domain due to the surge in publications, which makes keeping up with new findings challenging, necessitating automated knowledge extraction and processing. Of particular concern is the biomedical literature, where updates occur with ever-accelerating frequency (Fig. 1). Despite advances in healthcare, many diseases, such as Alzheimer’s disease (AD) (Scheltens et al., 2021; Trejo-Lopez et al., 2023) and multiple sclerosis (McGinley et al., 2021; Attfield et al., 2022), lack effective cures. Additionally, over 1,200 rare disorders have limited or no cures according to the National Organization for Rare Disorders. [^2] https://rarediseases.org/rare-diseases/ Discovering new scientific insights from research papers can expedite disease understanding and accelerate cure development.
@@
  • Two annotated datasets for RS and AD provide gold labels for semantic relations, aiding disease knowledge discovery research. [^3] The description of distantly supervised datasets for weakly supervised scenarios is included in Appendix H.
@@ 2 Data Pipeline
Mention extraction. MetaMapLite (Aronson, 2001; Demner-Fushman et al., 2017) (open-source BSD License) is provided by the National Library of Medicine (NLM) for extracting biomedical entities and mapping them to Concept Unique Identifiers (CUIs) within UMLS. The tool is updated every two years to incorporate the latest medical terminology and to ensure its accuracy in extraction and mapping. MetaMapLite simultaneously extracts mentions and links them to UMLS in one step, efficiently associating mentions with their corresponding CUIs. We detect a diverse range of entities, spanning 82 unique semantic types and covering a broad spectrum of biomedical concepts, including diseases, biologically active substances, anatomical structures, genes, and more. Detailed entity detection often leads to overlapping or successive entities in the text. To address this, our pipeline incorporates a merging strategy that consolidates overlapping or subsequent entities into cohesive units. For example, in the sentence: "To test norepinephrine augmentation as a potential disease-modifying therapy, we performed a biomarker-driven phase II trial of atomoxetine, a clinically-approved norepinephrine transporter inhibitor, in subjects with mild cognitive impairment due to AD.", the subsequent relevant mentions norepinephrine transporter and inhibitor are merged to one entity.
@@ 2.1 Dataset Creation
The result of the expert annotation yields two curated datasets. The Relation Detection dataset for Rett Syndrome (ReDReS) contains 601 sentences with 5,259 instances and 1,148 unique CUIs (Tab. 1). The inter-annotator agreement is measured using the Fleiss kappa score (McHugh, 2012), resulting in 0.6143 in the multi-class setup (4 classes) and indicating substantial agreement among annotators (Landis and Koch, 1977). In the binary setup (relation or no relation), the Fleiss kappa score is 0.7139. The Relation Detection dataset for Alzheimer’s Disease (ReDAD) comprises 641 sentences with 8,565 instances and 1,480 unique CUIs (Tab. 1). The Fleiss kappa score is 0.6403 in the multi-class setup and 0.7064 in the binary setup, showing substantial consensus among annotators. The final labels are determined through majority voting, leveraging the labels provided by each expert. Whereas the label distribution across classes is relatively balanced, the negative class is under-represented with 97 and 125 instances in ReDReS and ReDAD respectively (Tab. 1). Each dataset is randomly split into train, development, and test sets.
@@ 3 Models
Figure 3: Model Architecture of LaMReDA, LaMReDM (left), and LaMEL (right): Each model encodes the input sequence using BiomedBERT (large or base). For LaMReDA and LaMReDM, different tokens define the relation representation (A-P), passed through a linear projection layer, a dropout layer, and then a classification layer for prediction. The symbol # denotes element-wise addition and multiplication for LaMReDA and LaMReDM, respectively. For LaMEL, different tokens construct the entity representation (A-H), which are sent through a dropout layer and a linear layer to extract the projected entity representations.
@@ 3.1 LaMEL model
LaMEL learns an embedding space optimized for relation detection (Fig. 3). As the backbone language model (LM), we opt for BiomedBERT (Gu et al., 2021; Tinn et al., 2023) (MIT License), available in both uncased base and uncased large versions. [^8] HuggingFace’s Transformers library (Wolf et al., 2020) BiomedBERT is pretrained on the PubMed corpus, making it well-suited for our task as the curated datasets consist of sentences of abstracts from PubMed papers. Leveraging BiomedBERT ensures that the model can capture the language patterns prevalent in biomedical text. Following the LM encoding, we construct the representation of each entity by extracting its contextualized embedding $E_{i}$ corresponding to each entity $e_{i}$ from the encoded sequence. Subsequently, the entity representations are projected to the embedding space using a linear layer without changing the embedding dimension. The final prediction is based on cosine similarity between the two projected entity representations. If the cosine similarity exceeds a predefined threshold, the model predicts that there is a semantic relation between the two entities. We experiment with diverse strategies for learning entity representations (Fig. 3), aiming to optimize the effectiveness of the embedding space for the relation detection task. The explored types of entity representation $E$ are:
@@
where $\{E_{A},E_{B},E_{D},E_{E},E_{F},E_{G},E_{H}\}\in\mathbb{R}^{d}$ and $E_{C}\in\mathbb{R}^{2d}$, $d$ is the embedding size of BiomedBERT base (768) and BiomedBERT large (1024), ; defines the concatenation, * holds for the element-wise multiplication, $t_{[ent]}$, $t_{[/ent]}$ are the embeddings of the start and end special tokens of the entities, $[t_{E}]$ and $[t_{Inter}]$ are the averaged pooled representation of the entities and the intermediate tokens between the entities respectively.
@@ 3.2 LaMReD model
LaMReD provides two variations that differ in information synthesis (Fig. 3), aiming to explore the potential effect of different aggregations (Theodoropoulos and Moens, 2023). LaMReDA utilizes element-wise addition to aggregate the entities’ representations, whereas LaMReDM employs element-wise multiplication. The input text is encoded using BiomedBERT (base or large). Following LM encoding, we construct the relation representation by sampling and aggregating tokens from the input sequence. This step enables the model to capture essential features and contextual information relevant to semantic relation classification. To mitigate the risk of overfitting and enhance model generalization, we incorporate a dropout layer (Srivastava et al., 2014) with a probability of 0.3. The linear classification layer takes the aggregated representation and outputs the predicted label.
@@
where $\{R_{A},R_{B},R_{C},R_{D},R_{E},R_{F},R_{G},R_{H},R_{I},\\ R_{J},R_{K},R_{L},R_{M},R_{N},R_{O},R_{P}\}\in\mathbb{R}^{d}$, d is the embedding size of BiomedBERT base (768) and BiomedBERT large (1024), $f()$ is the aggregation function, element-wise addition for LaMReDA and element-wise multiplication for LaMReDM, $l()$ is a linear projection layer with dimension equal to the embedding size, $t_{[ent]_{1}}$, $t_{[/ent]_{1}}$, $t_{[ent]_{2}}$, and $t_{[/ent]_{2}}$ are the embeddings of the start and end special tokens of the first and second entity and $t_{[CLS]}$ is the representation of the special token [CLS]. We define the averaged pooled representation of the entities and the intermediate tokens between the entities as $[t_{E1}]$, $[t_{E2}]$, and $[t_{Inter}]$ correspondingly. In equations 23 and 24, we utilize the localized context vector $cv$ [^9] Additional information is provided in Appendix D. that utilizes the attention heads to locate relevant context for the entity pair and was introduced in ATLOP (Zhou et al., 2021), a state-of-the-art model in document-level relation extraction.
@@ 4 Results
Tables 2 and 3 report the F1-scores for LaMEL LaMReDA, and LaMReDM, models on the ReDReS and ReDAD datasets. Each cell (except for cross-disease experiments) displays two values: the average F1-score from 10 runs on the original test set (Tab. 1) and the average F1-score from a 5-fold cross-validation. The models perform well across all relation (A-P) and entity (A-H) representations, showing their ability to learn meaningful representations for the semantic relation task regardless of initial token selection. However, we observe patterns regarding the relation representations. In the binary setup, relation representation $R_{G}$ (Eq. 15) yields strong results for both datasets, suggesting that including the [CLS] token representation might be beneficial. In the multi-class setup, relation representations $R_{L}$ (Eq. 20), $R_{J}$ (Eq. 18), and $R_{O}$ (Eq. 23) are effective for both datasets, indicating that the surrounding context is crucial for the more complex task, as $R_{L}$ and $R_{J}$ include the averaged pooled representation of intermediate tokens between entities, and $R_{O}$ leverages the context vector (Zhou et al., 2021). The intra-model comparison reveals that over-parameterization tends to be useful. Using BiomedBERT large generally results in better performance than the base alternative. The BiomedBERT base shows superior performance mainly only in experiments using the original splits of ReDReS (Tab. 1). LaMEL is highly competitive with LaMReDA and LaMReDM, indicating that learning entity embedding spaces optimized for relation detection is promising. LaMEL achieves the highest performance in the 5-fold setup of ReDAD and the original setup of ReDReS, with F1-scores of 91.03% and 91.25% respectively.
@@
Table 2: LaMEL Results (%) in binary setup (BiomedBERT $\square$: base, $\blacksquare$:large): Each cell (unless cross-disease experiments) shows the average F1-score from 10 runs (original test set) and from 5-fold cross-validation setup.
The inter-model comparison across the same relation representations indicates that the aggregation function does not significantly impact relation detection tasks. Neither LaMReDA (element-wise addition) nor LaMReDM (element-wise multiplication) show a clear advantage over the other. This suggests that the transformer layers of BiomedBERT and the projection layer $l()$ preceding the aggregation are effectively trained in both models to encode the essential information for relation detection, regardless of the aggregation function used. The cross-disease experiments underscore the robustness of the models in both binary and multi-class setups. This robustness supports transfer learning (Zhuang et al., 2020) in semantic relation detection, extending to other diseases, highlighting the potential for broader applications and research endeavors in knowledge discovery.
@@
Table 3: LaMReDA and LaMReDM Results (%) in binary and multi-class setup (BiomedBERT $\square$:base,$\blacksquare$:large): Each cell (unless cross-disease experiments) shows the average F1-score from 10 runs (original test set) and from 5-fold cross-validation setup.
Human Performance. To assess and compare to human performance, two additional experts identify the relation type in a random sample of 300 instances from the test set of each dataset (Tab. 1). The evaluation ground truth is based on the original test set labels. In the binary setup, the average F1-score ranges from 92.14 for ReDReS to 91.87 for ReDAD. The LaMReDA, LaMReDM, and LaMEL models achieve performance comparable to human experts, indicating a high ability to detect semantic relations. Multi-class macro F1-scores range from 85.23 (micro: 85.45) to 85.76 (micro: 85.87) for ReDReS and ReDAD, respectively. Compared to human experts, all models show a performance gap, highlighting that identifying more complex aspects of semantic relation is a challenging task.
Baseline performance - lower bound. We randomly assign labels based on the training data’s class distribution (Tab 1). In the binary setup, the baseline achieves F1-scores of 54% (ReDReS) and 53.16% (ReDAD). For the multi-class setup, the macro F1-scores range from 32.05% to 32.43%, stressing the task’s difficulty, particularly for distinguishing various semantic relations (multi-class) [^11] More information is available in Appendix F..
@@ 5 Probing
This study probes BiomedBERT’s ability to capture semantic relations between entities. We explore different transformer layer representations and attention scores per layer and attention head. Averaged pooled entity representations are extracted from each layer, followed by training a linear classification layer. We test relation representations $R_{D}$, $R_{O}$, and $R_{P}$ (Eq. 12, 23, 24) of LaMReDA and LaMReDM to assess the impact of the context vector. Out-of-the-box representations are evaluated without the projection linear layer $l()$. We also extract average attention scores of tokens for each entity towards the other across each layer and head, concatenating these into a feature vector for training a linear classification layer. Following Chizhikova et al. (2022), we also train the classification layer using average attention scores between the two entities across all layers.
@@
Fig. 4 shows the results of probing experiments in the binary setup using ReDReS and BiomedBERT base. [^12] Additional probing experiments in Appendix I. The 10th and 11th layers provide the most informative representations for relation types ($R_{D}$, $R_{O}$, and $R_{P}$). The $R_{D}$ representation, using element-wise multiplication, outperforms other representations in intra-layer comparisons, suggesting its effectiveness without end-to-end training. However, as highlighted in section 4, inter-model comparisons indicate that the transformer layers and projection layer $l()$ capture crucial information for relation detection, regardless of the aggregation function. Using context vectors with $R_{O}$ and $R_{P}$ generally offers no advantage, though $R_{O}$ from the 10th and 11th layers performs well, indicating possibly meaningful localized context. Attention scores between entities in the 12th layer yield the best performance, surpassing the baseline of using scores from all layers, indicating strong attention between the entities in the last layer. Figure 4 reveals that the 6th and 9th attention heads are most informative for relation detection.